# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 5. Recognition and uniqueness results for groups and modules
Assume that $q=p^{m}$, where $p$ is odd, $G=\operatorname{SL}_{2}(q)$ and $V$ is a natural module for $G$ over $\mathbb{F}_{p}$. Set $H=\mathrm{GL}(V)\cong\mathrm{GL}_{2m}(p)$ and associate $G$ with its image in $H$. Let $S\in\mathrm{Syl}_{p}(G)$ and $B:=N_{G}(S)$. If $L\cong\operatorname{SL}_{2}(q)$ with $B<L\leq H$, then $L=G$.
@@
Let $L\cong\operatorname{SL}_{2}(q)$ be a subgroup of $H:=\mathrm{GL}_{2m}(p)$ containing $B:=N_{G}(S)$. Then $S\in\mathrm{Syl}_{p}(G)\cap\mathrm{Syl}_{p}(L)$ and $B=N_{L}(S)$. If $m=1$ then $H=\mathrm{GL}_{2}(p)$ has a unique subgroup isomorphic to $\operatorname{SL}_{2}(p)$ and so $L=G$. Note that the action of $S$ and Lemma 5.5 implies that there is $g\in H$ with $G^{g}=L$. We can arrange that $G\leq Y$ where $Y\cong\mathrm{GL}_{2}(q)$, and so $L\leq X\cong\mathrm{GL}_{2}(q)$ with $Y^{g}=X$. Now, $G=O^{p^{\prime}}(Y)$, $L=O^{p^{\prime}}(X)$, $Y=C_{H}(Z(Y))$ and $X=C_{H}(Z(X))$, and so to show that $G=L$, it suffices to prove that $Z(Y)=Z(X)$. Let $K$ be a complement to $S$ in $B$ so that $K$ is cyclic of order $q-1$. Then $Z(X)\leq C_{H}(K)\geq Z(Y)$.
@@
Let $V=V_{n}(p^{2})$ and $S=S_{n}(p^{2})$ for $11\leqslant n\leqslant p-$ 1, $G:=\mathrm{Aut}_{D^{\dagger}}(V)$ and $T:=\mathrm{Aut}_{S}(V)\in\mathrm{Syl}_{p}(\mathrm{Aut}_{D^{*}}(V))$. Suppose that $K\leq N_{G}(T)$ is abelian of order $(p-1)(p^{2}-1)$ such that $|C_{K}(T)|=p-1$. Assume further that there is $k\in K$ of order $p^{2}-1$ such that $k$ acts on $T$ as $\lambda^{-1}$ and on $C_{V}(T)$ as $\lambda$, where $\lambda\in\mathbb{F}_{p^{2}}^{*}$ has order $p^{2}-1$. Then $K=C_{K}(T)\times\langle k\rangle$, $KT$ is uniquely determined in $N_{\mathrm{Aut}_{D^{*}}(V)}(T)$ and $N_{\mathrm{Aut}(V)}(KT)=N_{G}(KT)$.
@@
[…]
By a Frattini argument, we have that $C_{N}(T)=TN_{C_{N}(T)}(K)$. Since $N_{C_{N}(T)}(K)$ normalises $T$ and $C_{V}(T)$ and as $\langle k\rangle$ is the unique subgroup of $K$ of order $q-1$ such that its element’s actions on $T$ and $C_{V}(T)$ are inverse to each other, we have that $N_{C_{N}(T)}(K)$ normalises $\langle k\rangle$. Then $[N_{C_{N}(T)}(K),\langle k\rangle]\leq C_{\langle k\rangle}(T)=1$. Moreover, $C_{K}(T)$ acts as scalars on $V$ and has order $p-1$ from which we deduce that $C_{K}(T)=Z(\mathrm{Aut}(V))$. Hence, $N_{C_{N}(T)}(K)$ centralises $K=C_{K}(T)\times\langle k\rangle$.
We have shown that $N=N_{\mathrm{Aut}_{D^{\dagger}}(V)}(KT)C_{N}(KT)$. Our aim will be to demonstrate that $C_{N}(KT)=Z(\mathrm{Aut}_{D^{*}}(V))$ has order $p^{2}-1$. Consider the restriction of $V$ to $\langle k\rangle$. Since $k$ acts as $\lambda^{-1}$ on $T$ and as $\lambda$ on $C_{V}(T)$, we have that $k$ corresponds to the element $\left(1,\lambda,\left(\begin{smallmatrix}1&0\\ 0&\lambda\end{smallmatrix}\right)\right)\in\mathrm{Aut}_{D^{*}}(V).$ Hence, $k$ acts as $\lambda^{n+1-i}$ on $[V,T;i]/[V,T;i+1]$ for $i\in\{0,\dots,n\}$. Since $10<n+1-i<p+$ 1, it follows that $k$ acts irreducibly on $[V,T;i]/[V,T;i+1]$ for all $i$. Moreover, since $\langle k\rangle$ has order $p^{2}-1$ and is centralised by $C_{N}(KT)$ we have that $C_{N}(KT)$ respects the $\mathbb{F}_{p^{2}}$-structure that $\langle k\rangle$ places on $V$.
@@
Let $q=p^{m}$ for $p$ an odd prime, $G$ be a finite group, $T\in\mathrm{Syl}_{p}(G)$ and $V$ be an elementary abelian $p$-group on which $G$ acts faithfully. Suppose that $L:=O^{p^{\prime}}(G)\cong\mathrm{(P)SL}_{2}(q)$ and $V|_{L}\cong V_{n}(q)|_{\mathbb{F}_{p}}$, where $12\leqslant n\leqslant p-$ 1, and recognise $G,\mathrm{Aut}_{D^{\dagger}}(V)\leq\mathrm{Aut}(V)$.
@@
Since $N_{L}(T)\leq N_{G\cap\mathrm{Aut}_{D}(V)}(\mathrm{Aut}_{S}(V))$, we recognise $T=\mathrm{Aut}_{S}(V)$. Let $X:=O^{p^{\prime}}(\mathrm{Aut}_{D}(V))$ and $K$ be an abelian subgroup of order greater than or equal to $\frac{(q-1)^{2}}{p+1}$ in a Hall $p^{\prime}$-subgroup of $N_{G\cap\mathrm{Aut}_{D}(V)}(T)$. Let $\widehat{K}$ be an abelian subgroup of order $(q-1)^{2}$ containing $K$ which lies in a Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{D}(V)}(T)$. Since $X\cong L\cong\mathrm{(P)SL}_{2}(q)$, we have that $C_{K}(T)=C_{K}(X)=C_{K}(L)$. Then $X,L\leq C_{\mathrm{Aut}(V)}(C_{K}(T))$. We have that $\mathrm{Aut}_{D}(V)$ preserves an $\mathbb{K}$-structure on $V$ and so $\mathrm{Aut}_{D}(V)\leq H\cong\mathrm{GL}_{n+1}(q)$ and $C_{K}(T)\leq Z(H)$ may be identified with a subgroup of the scalar matrices of $H$ acting on $V$. Since $H$ contains a Singer cycle and $|C_{K}(T)|\geqslant\frac{(q-1)^{2}}{p+1}$, we immediately see that $C_{\mathrm{Aut}(V)}(C_{K}(T))=H$ unless $m=2$ and $|C_{K}(T)|=p-1$. Thus, either $N_{\mathrm{Aut}(V)}(C_{K}(T))=N_{\mathrm{Aut}(V)}(H)$ or the conclusion of Lemma 5.18 are satisfied.
@@
Since $X\cong L\cong\mathrm{(P)SL}_{2}(q)$, we calculate that $N_{\mathrm{Aut}(V)}(X)=\mathrm{Aut}_{D^{\dagger}}(V)C_{\mathrm{Aut}(V)}(X)$. Observe that $|C_{\mathrm{Aut}_{D^{\dagger}}(V)}(X)|=q-1$ and that $C_{\mathrm{Aut}_{D^{\dagger}}(V)}(X)\leq C_{\mathrm{Aut}(V)}(X)$ acts transitively on vectors in $C_{V}(S)$. Let $Q$ be a $p$-subgroup of $C_{\mathrm{Aut}(V)}(X)$. Then $[Q,V]$ is normalised by $X$ and is properly contained in $V$ from which we conclude that $[Q,V]=1$ so that $Q=1$. Hence, $C_{\mathrm{Aut}(V)}(X)$ is a $p^{\prime}$-group. We deduce by Lemma 5.16 that $C_{\mathrm{Aut}(V)}(X)(K\cap N_{X}(T))=C_{C_{\mathrm{Aut}(V)}(X)(K\cap N_{X}(T))}(C_{V}(S))C_{\mathrm{Aut}_{D^{\dagger}}(V)}(X)$. Then, $C_{C_{\mathrm{Aut}(V)}(X)}(C_{V}(T))=1$ by the A×$\times$×B-Lemma [Asc86, (24.2)] and the claim holds.
@@
If $m=2$ and $|C_{K}(T)|=p-1$ then we observe that the conclusion of Lemma 5.18 holds and for $k\in K$ such that $k$ acts on $T$ as $\lambda^{-1}$ and on $C_{V}(T)$ as $\lambda$, where $\lambda\in\mathbb{F}_{p^{2}}^{*}$ has order $p^{2}-1$, the actions of $k^{ga}$ on $T/V$ and $C_{V}(T)$ are inverse to each other. Since $KT$ is uniquely determined in $N_{\mathrm{Aut}_{D}(V)}(\mathrm{Aut}_{S}(V)$ by Lemma 5.18, we deduce that $(KT)^{ga}=KT$. Since $N_{\mathrm{Aut}(V)}(KT)=N_{\mathrm{Aut}_{D^{\dagger}}(V)}(KT)$, we have that $ga$ normalises $X$ and as $X=L^{ga}$, we see that $X=L$ as desired.
@@
Now $J/T$ is an abelian group which acts faithfully and irreducibly on $S$. Hence $J/C_{J}(T)$ has order at most $q-1$ by Schur’s Lemma. Write $C_{J}(T)=T\times F$ for some abelian group $p^{\prime}$-subgroup $F$. Then $F$ acts on $C_{V}(T)$ and $F/C_{F}(C_{V}(T))$ embeds into $\mathrm{GL}_{1}(q)$ and so has order at most $q-1$. Now $T\times C_{F}(C_{V}(T))$ acts on $V$ and we deduce that $C_{F}(C_{V}(T))=1$ by the A×$\times$×B-Lemma [Asc86, (24.2)]. Hence, $|F|\leqslant q-1$, $|J/T|\leqslant(q-1)^{2}$ and the result holds. ∎
@@
Let $X:=O^{p^{\prime}}(\mathrm{Aut}_{D}(V))$ and $K$ be an abelian subgroup of order $(q-1)^{2}$ in a Hall $p^{\prime}$-subgroup of $N_{G\cap\mathrm{Aut}_{D}(V)}(T)$. We identify $T\in\mathrm{Syl}_{p}(G)$ with $\mathrm{Aut}_{S}(V)\leq\mathrm{Aut}_{D}(V)$. We have by Lemma 5.11 that $V|_{L}\cong\Lambda(q)|_{\mathbb{F}_{p}}\cong V|_{X}$. Set $W_{L},W_{X}$ to be the relevant $m(p-1)$-dimensional $\mathbb{F}_{p}$-submodules of $V$ for $L$ and $X$ respectively. We claim that $W_{L}=W_{X}$. Let $\lambda\in\mathbb{K}^{*}$ have order $q-1$, $r\in X$ be the conjugation map corresponding to $\left(1,1,\left(\begin{smallmatrix}\lambda^{-1}&0\\ 0&\lambda\end{smallmatrix}\right)\right)\in D,$ and $R=\langle r\rangle\leq X$ so that $|R|=q-1$. Then we may as well assume that $R\leq K$ and so normalises $X$ and $L$. Hence, it also normalises $W_{X}$ and $W_{L}$. Note that for $\overline{V}:=V/[V,T,T]$ we have that $\overline{[V,T]}=\overline{W_{X}}\times\overline{C_{V}(T)}=\overline{W_{L}}\times\overline{C_{V}(T)}$. Moreover, we calculate that $r$ acts as $\lambda^{-p}$ on $\overline{C_{V}(T)}\cong\overline{[V,T]}/\overline{W_{X}}$ and as $\lambda^{p-2}$ on $\overline{W_{X}}$. Since $q>p$, $\overline{C_{V}(T)}$ and $\overline{W_{X}}$ are non-isomorphic modules for $R$ and so $\overline{[V,T]}$ has precisely two proper non-trivial $R$-submodules. Since $\overline{W_{L}}\neq\overline{C_{V}(T)}$, we must have that $\overline{W_{X}}=\overline{W_{L}}$ so that $W:=W_{X}=W_{L}$, as claimed. Hence, $X,L\leq N_{\mathrm{Aut}(V)}(W)$ and we may consider the restriction of $N_{\mathrm{Aut}(V)}(W)$ to $\mathrm{Aut}(W)$. Note that $XK$ and $LK$ restrict faithfully to $W$ and applying Lemma 5.19, we deduce that $XC_{\mathrm{Aut}(V)}(W)=LC_{\mathrm{Aut}(V)}(W)$.
@@
Since $X\cong\operatorname{SL}_{2}(q)$ and $\mathrm{Aut}(X)=\mathrm{Aut}_{D^{\dagger}}(X)$, we calculate that $N_{\mathrm{Aut}(V)}(X)=\mathrm{Aut}_{D^{\dagger}}(V)C_{\mathrm{Aut}(V)}(X)$. Then $W$ is the unique proper non-trivial $X$-submodule of $V$ and we have that $C_{\mathrm{Aut}(V)}(X)$ normalises $W$. Indeed, since $XC_{\mathrm{Aut}(V)}(W)=LC_{\mathrm{Aut}(V)}(W)$, it follows from the proof of Lemma 5.19 that $N_{\mathrm{Aut}(V)}(X)=\mathrm{Aut}_{D^{\dagger}}(V)C_{C_{\mathrm{Aut}(V)}(X)}(W)$. Let $Q$ be a $p$-subgroup of $C_{C_{\mathrm{Aut}(V)}(X)}(W)$. Then $[Q,V]$ and $[Q,W]$ are normalised by $X$ and properly contained in $V$, resp. $W$, from which we conclude that $Q\leq C_{C_{\mathrm{Aut}(V)}(X)}(W)\cap C_{C_{\mathrm{Aut}(V)}(X)}(V/W)$. Then $C_{\mathrm{Aut}_{D^{\dagger}}(V)}C_{C_{\mathrm{Aut}(V)}(X)}(W)/C_{\mathrm{Aut}(V)}(V/W)$ is a $p^{\prime}$-group acting transitively on vectors in $C_{V/W}(T)$. We conclude by Lemma 5.16 and the A×$\times$×B-Lemma [Asc86, (24.2)] that
@@
Aiming for a contradiction, assume that $H\neq 1$ so that $[H,V]=W$. Note that $[[V,T],T,H]=1=[T,H,[V,T]]$ from which we conclude that $[[V,T],H,T]=1$ and since $[V,H]\leq W$, we must have that $[[V,T],H]\leq C_{W}(T)$. On the other hand, $[T,H,V]=1$ and we deduce that $[H,V,T]=[W,T]\leq C_{W}(T)$. Since $p>3$, this is a contradiction.
@@
On the other hand, $T^{g}\in\mathrm{Syl}_{p}(X)$ and since $c:=c_{1}c_{2}\in N_{\mathrm{Aut}(V)}(X)$, we deduce that $T^{gc}\in\mathrm{Syl}_{p}(X)$. Hence there is $a\in X$ such that $T^{gca}=T$. Since $K^{gc}$ normalises $T^{gc}$, we ascertain that $K^{gca}\leq N_{\mathrm{Aut}_{D^{\dagger}}(V)}(T)$ and since $K^{gca}$ is elementary abelian of order $(q-1)^{2}$, we conclude that $(TK)^{gca}=TK^{gca}=TK$. Finally, we have that $L^{gca}=X^{ca}=X$. Setting $B=TK$, and replacing $gca$ by $g$, we have demonstrated that there is $g\in N_{\mathrm{Aut}(V)}(B)$ with $L^{g}=X$. To complete the proof of the lemma, we shall show that $N_{\mathrm{Aut}(V)}(B)$ normalises $X$.
@@
From the proof of Lemma 5.19, we have that $N_{\mathrm{Aut}(V)}(B)\leq C_{\mathrm{Aut}(V)}(W)N_{\mathrm{Aut}_{D^{*}}(V)}(B)$. It remains to show that $N_{C_{\mathrm{Aut}(V)}(W)}(B)=1$. Indeed, we have that $N_{C_{\mathrm{Aut}(V)}(W)}(B)$ normalises $T$ and $N_{C_{\mathrm{Aut}(V)}(W)}(B)=N_{N_{C_{\mathrm{Aut}(V)}(W)}(B)}(K)$ by the Frattini argument. Suppose that $\ell\in N_{C_{\mathrm{Aut}(V)}(W)}(T)$ has $p^{\prime}$-order. Then $[\ell,T]=1$ and $V=[V,\ell]\times C_{V}(\ell)$ is a $T$-invariant decomposition. Since $\ell\in C_{\mathrm{Aut}(V)}(W)$, we have that $W\leq C_{V}(\ell)$ and $C_{V}(\ell)\neq 1$. Observe that for $U\in\{[V,\ell],C_{V}(\ell)\}$, if $U\not\leq[V,T]$ then $[V,T]=[U,T][V,T,T]$. Since $T$ acts nilpotently on $V$, we deduce that $C_{V}(T)\leq[V,T]\leq U$ and since $V=[V,\ell]\times C_{V}(\ell)$ is a $T$-invariant decomposition, we conclude that $U=V$. Finally, as $C_{V}(\ell)\neq 1$, we deduce that $V=C_{V}(\ell)$ and $\ell=1$. Therefore, $N_{C_{\mathrm{Aut}(V)}(W)}(B)$ is a $p$-group.

### v2

@@ 5. Recognition and uniqueness results for groups and modules
Assume that $q=p^{m}$, where $p$ is odd, $G=\operatorname{SL}_{2}(q)$ and $V$ is a natural module for $G$ over $\mathbb{F}_{p}$. Set $H=\mathrm{GL}(V)\cong\mathrm{GL}_{2m}(p)$ and associate $G$ with its image in $H$. Let $S\in\mathrm{Syl}_{p}(G)$ and $N_{G}(S)$. If $L\cong\operatorname{SL}_{2}(q)$ with $N_{G}(S)<L\leq H$, then $L=G$.
@@
Let $L\cong\operatorname{SL}_{2}(q)$ be a subgroup of $H:=\mathrm{GL}_{2m}(p)$ containing $N_{G}(S)$. Then $S\in\mathrm{Syl}_{p}(G)\cap\mathrm{Syl}_{p}(L)$ and $N_{G}(S)=N_{L}(S)$. If $m=1$ then $H=\mathrm{GL}_{2}(p)$ has a unique subgroup isomorphic to $\operatorname{SL}_{2}(p)$ and so $L=G$. Note that the action of $S$ and Lemma 5.5 implies that there is $g\in H$ with $G^{g}=L$. We can arrange that $G\leq Y$ where $Y\cong\mathrm{GL}_{2}(q)$, and so $L\leq X\cong\mathrm{GL}_{2}(q)$ with $Y^{g}=X$. Now, $G=O^{p^{\prime}}(Y)$, $L=O^{p^{\prime}}(X)$, $Y=C_{H}(Z(Y))$ and $X=C_{H}(Z(X))$, and so to show that $G=L$, it suffices to prove that $Z(Y)=Z(X)$. Let $K$ be a complement to $S$ in $N_{G}(S)$ so that $K$ is cyclic of order q1q-1. Then $Z(X)\leq C_{H}(K)\geq Z(Y)$.
@@
Recall the groups $D$, $D^{*}$ and $D^{\dagger}$ from Notation 2.1 and the groups $P_{n}(q)$ and $P_{n}^{*}(q)$ from Notation 2.3.
@@
Let $V=V_{n}(p^{2})$ and $S=S_{n}(p^{2})$ for $1\leqslant n\leqslant p-1$, $G:=\mathrm{Aut}_{D^{\dagger}}(V)$ and $T:=\mathrm{Aut}_{S}(V)\in\mathrm{Syl}_{p}(\mathrm{Aut}_{D^{*}}(V))$. Suppose that $K\leq N_{G}(T)$ is abelian of order $(p-1)(p^{2}-1)$ such that $|C_{K}(T)|=p-1$. Assume further that there is $k\in K$ of order $p^{2}-1$ such that $k$ acts on $T$ as $\lambda^{-1}$ and on $C_{V}(T)$ as $\lambda$, where $\lambda\in\mathbb{F}_{p^{2}}^{*}$ has order $p^{2}-1$. Then $K=C_{K}(T)\times\langle k\rangle$, $KT$ is uniquely determined in $N_{\mathrm{Aut}_{D^{*}}(V)}(T)$ and $N_{\mathrm{Aut}(V)}(KT)=N_{G}(KT)$.
@@
Set $N:=N_{\mathrm{Aut}(V)}(KT)$ so that $T\trianglelefteq N$. Since $T$ is elementary abelian of order $p^{2}$, we have that $N/C_{N}(T)$ is isomorphic to a subgroup of $\mathrm{GL}_{2}(p)$. Moreover, $KC_{N}(T)/C_{N}(T)$ has order $p^{2}-1$ and is normal in $N/C_{N}(T)$ from which we deduce that $|N/C_{N}(T)|\leqslant(p^{2}-1).2$. Since $|N_{\mathrm{Aut}_{D^{\dagger}}(V)}(KT)/C_{\mathrm{Aut}_{D^{\dagger}}(V)}(KT)|=(p^{2}-1).2$, we have that $N=N_{\mathrm{Aut}_{D^{\dagger}}(V)}(KT)C_{N}(T)$.
By a Frattini argument, we have that $C_{N}(T)=TN_{C_{N}(T)}(K)$. Since $N_{C_{N}(T)}(K)$ normalises $T$ and $C_{V}(T)$ and as $\langle k\rangle$ is the unique subgroup of $K$ of order q1q-1 such that its element’s actions on $T$ and $C_{V}(T)$ are inverse to each other, we have that $N_{C_{N}(T)}(K)$ normalises $\langle k\rangle$. Then $N_{C_{N}(T)}(K),\langle k\rangle]\leq C_{\langle k\rangle}(T)=1$. Moreover, $C_{K}(T)$ acts as scalars on $V$ and has order p1p-1 from which we deduce that $C_{K}(T)=Z(\mathrm{Aut}(V))$. Hence, $N_{C_{N}(T)}(K)$ centralises $K=C_{K}(T)\times\langle k\rangle$.
We have shown that $N=N_{\mathrm{Aut}_{D^{\dagger}}(V)}(KT)C_{N}(KT)$. Our aim will be to demonstrate that $C_{N}(KT)=Z(\mathrm{Aut}_{D^{*}}(V))$ has order $p^{2}-1$. Consider the restriction of $V$ to $\langle k\rangle$. Since $k$ acts as $\lambda^{-1}$ on $T$ and as $\lambda$ on $C_{V}(T)$, we have that $k$ corresponds to the element $\left(1,\lambda,\left(\begin{smallmatrix}1&0\\ 0&\lambda\end{smallmatrix}\right)\right)\in\mathrm{Aut}_{D^{*}}(V).$ Hence, $k$ acts as $\lambda^{n+1-i}$ on $[V,T;i]/[V,T;i+1]$ for $i\in\{0,\dots,n\}$. Since 0<n+1i<p+10<n+1-i<p+1, it follows that $k$ acts irreducibly on $[V,T;i]/[V,T;i+1]$ for all $i$. Moreover, since $\langle k\rangle$ has order $p^{2}-1$ and is centralised by $C_{N}(KT)$ we have that $C_{N}(KT)$ respects the $\mathbb{F}_{p^{2}}$-structure that $\langle k\rangle$ places on $V$.
@@
Let $q=p^{m}$ for $p$ an odd prime, $G$ be a finite group, $T\in\mathrm{Syl}_{p}(G)$ and $V$ be an elementary abelian $p$-group on which $G$ acts faithfully. Suppose that $L:=O^{p^{\prime}}(G)\cong\mathrm{(P)SL}_{2}(q)$ and $V|_{L}\cong V_{n}(q)|_{\mathbb{F}_{p}}$, where $2\leqslant n\leqslant p-1$, and recognise $G,\mathrm{Aut}_{D^{\dagger}}(V)\leq\mathrm{Aut}(V)$.
@@
Since $N_{L}(T)\leq N_{G\cap\mathrm{Aut}_{D}(V)}(\mathrm{Aut}_{S}(V))$, we recognise $T=\mathrm{Aut}_{S}(V)$. Let $X:=O^{p^{\prime}}(\mathrm{Aut}_{D}(V))$ and $K$ be an abelian subgroup of order greater than or equal to $\frac{(q-1)^{2}}{p+1}$ in a Hall $p^{\prime}$-subgroup of $N_{G\cap\mathrm{Aut}_{D}(V)}(T)$. Let $\widehat{K}$ be an abelian subgroup of order $(q-1)^{2}$ containing $K$ which lies in a Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{D}(V)}(T)$. Since $X\cong L\cong\mathrm{(P)SL}_{2}(q)$, we have that $C_{K}(T)=C_{K}(X)=C_{K}(L)$. Then $X,L\leq C_{\mathrm{Aut}(V)}(C_{K}(T))$. We have that $\mathrm{Aut}_{D}(V)$ preserves an $\mathbb{K}$-structure on $V$, where $\mathbb{K}:=\mathbb{F}_{q}$. Thus, $\mathrm{Aut}_{D}(V)\leq H\cong\mathrm{GL}_{n+1}(q)$ and $C_{K}(T)\leq Z(H)$ may be identified with a subgroup of the scalar matrices of $H$ acting on $V$. Since $H$ contains a Singer cycle and $|C_{K}(T)|\geqslant\frac{(q-1)^{2}}{p+1}$, we immediately see that $C_{\mathrm{Aut}(V)}(C_{K}(T))=H$ unless $m=2$ and $|C_{K}(T)|=p-1$. Thus, either $N_{\mathrm{Aut}(V)}(C_{K}(T))=N_{\mathrm{Aut}(V)}(H)$ or the conclusions of Lemma 5.18 are satisfied.
@@
[…]
@@
If $m=2$ and $|C_{K}(T)|=p-1$ then we observe that the conclusion of Lemma 5.18 holds and for $k\in K$ such that $k$ acts on $T$ as $\lambda^{-1}$ and on $C_{V}(T)$ as $\lambda$, where $\lambda\in\mathbb{F}_{p^{2}}^{*}$ has order $p^{2}-1$, the actions of $k^{ga}$ on $T/V$ and $C_{V}(T)$ are inverse to each other. Since $KT$ is uniquely determined in $N_{\mathrm{Aut}_{D}(V)}(\mathrm{Aut}_{S}(V))$ by Lemma 5.18, we deduce that $(KT)^{ga}=KT$. Since $N_{\mathrm{Aut}(V)}(KT)=N_{\mathrm{Aut}_{D^{\dagger}}(V)}(KT)$, we have that $ga$ normalises $X$ and as $X=L^{ga}$, we see that $X=L$ as desired.
@@
[…]
@@
Let $X:=O^{p^{\prime}}(\mathrm{Aut}_{D}(V))$ and $K$ be an abelian subgroup of order $(q-1)^{2}$ in a Hall $p^{\prime}$-subgroup of $N_{G\cap\mathrm{Aut}_{D}(V)}(T)$. We identify $T\in\mathrm{Syl}_{p}(G)$ with $\mathrm{Aut}_{S}(V)\leq\mathrm{Aut}_{D}(V)$. We have by Lemma 5.11 that $V|_{L}\cong\Lambda(q)|_{\mathbb{F}_{p}}\cong V|_{X}$. Set $W_{L},W_{X}$ to be the relevant m(p1)m(p-1)-dimensional $\mathbb{F}_{p}$-submodules of $V$ for $L$ and $X$ respectively. We claim that $W_{L}=W_{X}$. Let $\lambda\in\mathbb{K}^{*}$ have order q1q-1, $r\in X$ be the conjugation map corresponding to $\left(1,1,\left(\begin{smallmatrix}\lambda^{-1}&0\\ 0&\lambda\end{smallmatrix}\right)\right)\in D,$ and $R=\langle r\rangle\leq X$ so that |R|=q1|R|=q-1. Then we may as well assume that $R\leq K$ and so normalises $X$ and $L$. Hence, it also normalises $W_{X}$ and $W_{L}$. Note that for $\overline{V}:=V/[V,T,T]$ we have that $\overline{[V,T]}=\overline{W_{X}}\times\overline{C_{V}(T)}=\overline{W_{L}}\times\overline{C_{V}(T)}$. Moreover, we calculate that $r$ acts as $\lambda^{-p}$ on $\overline{C_{V}(T)}\cong\overline{[V,T]}/\overline{W_{X}}$ and as $\lambda^{p-2}$ on $\overline{W_{X}}$. Since $q>p$, $\overline{C_{V}(T)}$ and $\overline{W_{X}}$ are non-isomorphic modules for $R$ and so $\overline{[V,T]}$ has precisely two proper non-trivial $R$-submodules. Since $\overline{W_{L}}\neq\overline{C_{V}(T)}$, we must have that $\overline{W_{X}}=\overline{W_{L}}$ so that $W:=W_{X}=W_{L}$, as claimed. Hence, $X,L\leq N_{\mathrm{Aut}(V)}(W)$ and we may consider the restriction of $N_{\mathrm{Aut}(V)}(W)$ to $\mathrm{Aut}(W)$. Note that $XK$ and $LK$ restrict faithfully to $W$ and applying Lemma 5.19, we deduce that $XC_{\mathrm{Aut}(V)}(W)=LC_{\mathrm{Aut}(V)}(W)$.
@@
[…]
@@
Aiming for a contradiction, assume that $H\neq 1$ so that $[H,V]=W$. Note that $[[V,T],T,H]=1=[T,H,[V,T]]$ from which we conclude that $[[V,T],H,T]=1$ and since $V,H]\leq W$, we must have that $V,T],H]\leq C_{W}(T)$. On the other hand, $[T,H,V]=1$ and we deduce that $H,V,T]=[W,T]\leq C_{W}(T)$. Since $p>3$, this is a contradiction.
@@
On the other hand, $T^{g}\in\mathrm{Syl}_{p}(X)$ and since $c:=c_{1}c_{2}\in N_{\mathrm{Aut}(V)}(X)$, we deduce that $T^{gc}\in\mathrm{Syl}_{p}(X)$. Hence there is $a\in X$ such that $T^{gca}=T$. Since $K^{gc}$ normalises $T^{gc}$, we ascertain that $K^{gca}\leq N_{\mathrm{Aut}_{D^{\dagger}}(V)}(T)$ and since $K^{gca}$ is elementary abelian of order $(q-1)^{2}$, we conclude that $(TK)^{gca}=TK^{gca}=TK$. Finally, we have that $L^{gca}=X^{ca}=X$. Replacing gcagca by $g$, we have demonstrated that there is $g\in N_{\mathrm{Aut}(V)}(TK)$ with $L^{g}=X$. To complete the proof of the lemma, we shall show that $N_{\mathrm{Aut}(V)}(TK)$ normalises $X$.
@@
From the proof of Lemma 5.19, we have that $N_{\mathrm{Aut}(V)}(TK)\leq C_{\mathrm{Aut}(V)}(W)N_{\mathrm{Aut}_{D^{*}}(V)}(TK)$. It remains to show that $N_{C_{\mathrm{Aut}(V)}(W)}(B)=1$. Indeed, we have that $N_{C_{\mathrm{Aut}(V)}(W)}(B)$ normalises $T$ and $N_{C_{\mathrm{Aut}(V)}(W)}(B)=N_{N_{C_{\mathrm{Aut}(V)}(W)}(TK)}(K)$ by the Frattini argument. Suppose that $\ell\in N_{C_{\mathrm{Aut}(V)}(W)}(T)$ has $p^{\prime}$-order. Then $\ell,T]=1$ and $V=[V,\ell]\times C_{V}(\ell)$ is a $T$-invariant decomposition. Since $\ell\in C_{\mathrm{Aut}(V)}(W)$, we have that $W\leq C_{V}(\ell)$ and $C_{V}(\ell)\neq 1$. Observe that for $U\in\{[V,\ell],C_{V}(\ell)\}$, if $U\not\leq[V,T]$ then $[V,T]=[U,T][V,T,T]$. Since $T$ acts nilpotently on $V$, we deduce that $C_{V}(T)\leq[V,T]\leq U$ and since $V=[V,\ell]\times C_{V}(\ell)$ is a $T$-invariant decomposition, we conclude that $U=V$. Finally, as $C_{V}(\ell)\neq 1$, we deduce that $V=C_{V}(\ell)$ and $\ell=1$. Therefore, $N_{C_{\mathrm{Aut}(V)}(W)}(TK)$ is a $p$-group.